# Linear-scaling Tensor Train Sketching

### 11 Mar 2026 v1

@@ 4.2 TT Random Projections
From Definition 4.2 it is clear that $f_{\mathtt{TT}(R)}$ can be seen as a down-sampled Block-Sparse Tensor Train sketch. The exponential scaling requirement in $P$ may thus be quite pessimistic, with conditions similar to those of Theorems 3.7 and 3.10 being sufficient, but we do not pursue this analysis in this work.
@@ 5.1 Algorithms
5.2 Application of the Block-Sparse Tensor Train sketch
The key insight in the application of the Gaussian TT sketch – and by extension of BSTT – is its natural implementation as a recursive sequence of contractions. Given ${\bm{\Omega}}_{\mathtt{GTT}}=(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}$, with uniform ranks $R$ for simplicity, and the vector $\mathbf{x}$ with TT representation $\mathcal{X}=\mathcal{X}_{1}\bowtie\cdots\bowtie\mathcal{X}_{d}$ and ranks $s_{1},\dots,s_{d-1}$, we can formulate ${\bm{\Omega}}_{\mathtt{GTT}}\mathbf{x}$ in terms of matrix unfoldings as follows:
@@ 5.2 Application of the Block-Sparse Tensor Train sketch
The corresponding BSTT sketch is then obtained by vertically stacking independent realizations:
@@
This structure inherently enables parallel execution and should seamlessly extend to adaptive algorithms where target ranks are not known a priori and the sketch is incrementally enlarged, analogous to the Khatri-Rao formulation presented in [12], although this is left for future work. Each intermediate matrix $\mathbf{W}_{k}$ for $k=2,\dots,d$ is itself a BSTT sketch of the partial block tensor train obtained by contracting only cores to the right of mode $k$:
@@
using the notation from Remark 3.3. Therefore the computation captures, at no additional cost, the row space of each partial unfolding $\mathbf{X}_{k}^{>}:=(\mathcal{X}_{k}\bowtie\cdots\bowtie\mathcal{X}_{d})^{\leq 1}$ for $k\in[d]$ in addition to ${\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{x}$ being sketched using matrices ${\bm{\Omega}}^{>}_{\mathtt{BSTT},k}:=(\mathcal{G}_{k}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}\in\mathbb{F}^{PR\times n_{k}\dots n_{d}}$. We note that these matrices are themselves BSTT sketches and inherit their theoretical guarantees (Theorems 3.7 and 3.10), although they are clearly not independent of one another.
@@
Algorithm 4 Partial Contractions with Block-Sparse Tensor Train
@@
2:Partial sketches $(\mathbf{W}_{1},\dots,\mathbf{W}_{d})$ with $\mathbf{W}_{k}={\bm{\Omega}}^{>}_{\mathtt{BSTT},k}(\mathbf{X}^{>}_{k})^{T}\in\mathbb{F}^{Pr_{k-1}\times r_{k-1}}$.
@@ 5.2.1 Linear combinations
Explicitly assembling these block-sparse cores of size $J\chi\times n\times J\chi$ and then applying Algorithm 4 is clearly an ineffective way to compute the partial contractions. Indeed, it can be readily seen [4] that the partial sketches $\mathbf{W}_{k}={\bm{\Omega}}^{>}_{\mathtt{BSTT},k}(\mathbf{S}_{k}^{>})^{T}$ for $k\geq 2$ can be formed by vertically stacking the independent partial sketches $\mathbf{W}_{k,j}={\bm{\Omega}}^{>}_{\mathtt{BSTT},k}(\mathbf{S}^{>}_{k,j})^{T}$ for $j\in[J]$, obtained by applying Algorithm 4 independently to each summand TT with the sketch cores:
$\mathbf{W}_{k}=\begin{bmatrix}\alpha_{1}\mathbf{W}_{k,1}\\ \vdots\\ \alpha_{j}\mathbf{W}_{k,J}\end{bmatrix}\text{ for }k=d,\dots,2\quad\text{ and }\quad{\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{s}=\mathbf{w}_{1}=\sum_{j=1}^{J}\alpha_{j}\mathbf{w}_{1,j}=\sum_{j=1}^{J}\alpha_{j}{\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{y}_{j},$
@@ 5.2.2 Matrix-vector product
In the first step, the contraction costs $\mathcal{O}(nPRr_{H}(\chi+n))$ algebraic operations. The following one costs $\mathcal{O}(nPRr_{H}\chi(nr_{H}+R+\chi))$ algebraic operations. Proceeding recursively, the tensor contraction of ${\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{H}\mathbf{y}$ costs in total $\mathcal{O}(dnPRr_{H}\chi(nr_{H}+R+\chi))$ algebraic operations. This is significantly better than assembling the tensor train of ranks $R_{H}\chi$ representing $\mathbf{H}\mathbf{y}$ at a cost of $\mathcal{O}(d\chi^{2}n^{2}r_{H}^{2})$, then applying Algorithm 4 with a cost of $\mathcal{O}\bigl(dnPRr_{H}\chi(R+r_{H}\chi)\bigr)$.
(a) Tensor network representation of ${\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{H}\mathbf{y}$
@@
Figure 4: Graphical representation of the sketching of the matrix-vector product $\mathbf{H}\mathbf{y}$. The matrix $\mathbf{H}$ has a TT representation $(\mathcal{H}_{k})_{k\in[d]}$ of ranks $r_{H}$, $\mathbf{y}$ has a TT representation $(\mathcal{Y}_{k})_{k\in[d]}$ of ranks $\chi$, ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$. The contractions are done from right to left, top to bottom.
@@ 5.3 Numerical Experiments
In this section, we present numerical tests of the proposed BSTT sketch first in synthetic examples, then in the context of TT-rounding for compressing Hadamard products of tensors, and finally showcase a tensor train-adapted sketched Rayleigh-Ritz eigensolver applied to an example from quantum chemistry.
@@ 5.3.1 A Synthetic Example
We apply both Algorithms 2 and 3 with orthogonal Block-Sparse Tensor Train sketches to compress a synthetic perturbed low-rank tensor $\mathcal{X}=\mathcal{X}_{s}+\varepsilon\mathcal{X}_{10}$, where $\mathcal{X}_{10}$ is a rank-10 Gaussian TT perturbation with variable noise level $\varepsilon=10^{-s}$. We consider two structures for the dominant term $\mathcal{X}_{s}$: a normalized sum of 16 rank-1 Gaussian Kronecker TT vectors, and a single rank-16 Gaussian TT vector. In both cases we fix the embedding dimension $PR=16$ and vary the block rank $R\in\{1,4,8,16\}$. By Theorem 3.10, increasing $R$ at fixed $PR$ reduces the subspace coherence constant $C_{\mathbf{Q}}(R)$, improving the injectivity parameter and, via Theorem 3.14, the rounding error bound. The two choices of $\mathcal{X}_{s}$ probe how the TT-rank structure of the dominant component interacts with this block-rank dependence.
@@ 5.3.2 Hadamard Product
Despite this representational efficiency, performing pointwise algebraic operations within the QTT format remains a significant computational bottleneck. We apply Algorithm 2 with orthogonal BSTT cores to compress a three-term Hadamard product, and report the accuracy and computation time as functions of the target rank.
Figure 7: Randomized rounding of Hadamard products of QTT functions. Empirical performance of Orthogonal BSTT sketches in the randomized rounding of a three-term QTT Hadamard product. The sketches are applied via Algorithm 2 to the functions defined in (11), evaluated on the discretized domain $0,1]^{3}$ over a $(2^{20})^{3}$ grid. We compare embedding dimensions $PR=r$ across varying block ranks $R\in\{1,4,8,16\}$, where markers indicate the median across 1000 independent trials, and shaded regions denote the interquartile range (25th to 75th percentiles).
@@ 5.3.3 Ground-state energy of a quantum chemistry Hamiltonian
In this experiment, we apply the Orthogonal Block-Sparse Tensor Train sketch to the computation of the electronic ground-state energy of the lithium hydride (LiH) molecule in the ccpvdz basis.
The electronic Hamiltonian is written in second quantization [46], where it is naturally represented as a TTO. For the LiH molecule, the Hamiltonian corresponds to a symmetric matrix of size $2^{38}\times 2^{38}$. Starting from a Hartree-Fock initial guess $\mathcal{V}_{0}$, which has TT rank 1, we construct a Krylov basis by repeatedly applying the TTO and rounding the result via Algorithm 2 to maintain a bounded TT rank. The Orhogonal BSTT sketch plays the dual role of providing the random projections needed for randomized rounding of the new basis vector $\mathbf{b}_{j}=\mathbf{H}\mathbf{b}_{j-1}-\sum_{s<j}\alpha_{s}\mathbf{b}_{s}$ at each step, and also enabling a sketched Rayleigh-Ritz approach following [33].
@@
5:  Draw sketch ${\bm{\Omega}}_{\mathtt{BSTT}}=(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}\in\ \mathbb{F}^{PR\times N}$ with parameters $(P,R)$
@@
Given the sketched basis matrix $\mathbf{C}=[{\bm{\Omega}}_{OBSTT}\,\mathbf{b}_{1}\cdots{\bm{\Omega}}_{OBSTT}\,\mathbf{b}_{m}]$ and the sketched action $\mathbf{D}=[{\bm{\Omega}}_{OBSTT}\,\mathbf{H}\mathbf{b}_{1}\cdots{\bm{\Omega}}_{OBSTT}\,\mathbf{H}\mathbf{b}_{m}]$, the Ritz matrix is obtained as the least-squares solution $\mathbf{M}=\mathbf{C}^{\dagger}\mathbf{D}$ via a QR decomposition of $\mathbf{C}$, and then diagonalized to yield the Ritz pairs and in particular an approximation $\lambda_{1}$ to the lowest eigenvalue of $\mathbf{H}$ and corresponding approximate eigenvector $\widehat{\mathbf{x}}=\sum_{j=1}^{m}y_{j}\mathbf{b}_{j}$ with $\mathbf{M}\mathbf{y}=\lambda_{1}\mathbf{y}$. In order to control the conditioning of the basis, we additionally implement a sketched Gram-Schmidt algorithm, meaning that our basis vectors are (partially) sketch-orthogonal, $({\bm{\Omega}}_{OBSTT}\mathbf{b}_{i})^{*}({\bm{\Omega}}_{OBSTT}\mathbf{b}_{j})\approx\delta_{ij}$, with inaccurate evaluations due to the TT-rounding steps necessary to keep ranks under control. This ensures good conditioning of the basis when ${\bm{\Omega}}_{OBSTT}$ is an OSE for the (approximate) Krylov subspace. The same sketch is reused to (1) round matrix-vector products and linear combinations, (2) compute the coefficients needed to sketch-orthogonalize the Krylov basis, and (3) form the sketched problem. The resulting algorithm is summarized in Algorithm 5.
@@
Convergence results are reported on Figure 8. For each quantity of interest, namely eigenvalue error and eigenvector residual, we compare the quantities obtained from the sketched problem at each iteration, $\lambda_{1}-E_{\mathrm{ref}}$ and $\|\mathbf{D}\mathbf{y}-\lambda_{1}\mathbf{C}\mathbf{y}\|^{2}_{2}$, and the true residuals $\langle\widehat{\mathbf{x}}|\mathbf{H}|\widehat{\mathbf{x}}\rangle-E_{ref}$ and $\|\mathbf{H}\widehat{\mathbf{x}}-\lambda_{1}\widehat{\mathbf{x}}\|^{2}_{2}$. Our observation is that the approximate eigenvalue $\lambda_{1}$ is quite noisy and inaccurate, however evaluation of the true Rayleigh quotient shows a fairly steady convergence behavior, with 5 digits of accuracy obtained within 80 iterations. On the other hand, the sketched residual remains a good approximation of the true residual, as expected from the OSE / OSI guarantees of the BSTT sketch.
6 Oblivious Subspace Embedding (OSE) property for Block-Sparse Tensor Train
This section begins the technical development of this paper, where we establish OSE and OSI guarantees by rigorous probabilistic techniques. To provide a pedagogical exposition, we provide full statements for the key results needed for the analysis of tensor-formatted sketch matrices such as ${\bm{\Omega}}_{\mathtt{GTT}}$ and ${\bm{\Omega}}_{\mathtt{BSTT}}$.

### 13 Mar 2026 v2

@@ 4.2 TT Random Projections
From Definition 4.2 it is clear that $f_{\mathtt{TT}(R)}$ can be seen as a down-sampled TTStack sketch. The exponential scaling requirement in $P$ may thus be quite pessimistic, with conditions similar to those of Theorems 3.7 and 3.10 being sufficient, but we do not pursue this analysis in this work.
@@ 5.1 Algorithms
5.2 Application of the TTStack sketch
The key insight in the application of the Gaussian TT sketch – and by extension of TTStack – is its natural implementation as a recursive sequence of contractions. Given ${\bm{\Omega}}_{\mathtt{GTT}}=(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}$, with uniform ranks $R$ for simplicity, and the vector $\mathbf{x}$ with TT representation $\mathcal{X}=\mathcal{X}_{1}\bowtie\cdots\bowtie\mathcal{X}_{d}$ and ranks $s_{1},\dots,s_{d-1}$, we can formulate ${\bm{\Omega}}_{\mathtt{GTT}}\mathbf{x}$ in terms of matrix unfoldings as follows:
@@ 5.2 Application of the TTStack sketch
The corresponding TTStack sketch is then obtained by vertically stacking independent realizations:
@@
This structure inherently enables parallel execution and should seamlessly extend to adaptive algorithms where target ranks are not known a priori and the sketch is incrementally enlarged, analogous to the Khatri-Rao formulation presented in [12], although this is left for future work. Each intermediate matrix $\mathbf{W}_{k}$ for $k=2,\dots,d$ is itself a TTStack sketch of the partial block tensor train obtained by contracting only cores to the right of mode $k$:
@@
using the notation from Remark 3.3. Therefore the computation captures, at no additional cost, the row space of each partial unfolding $\mathbf{X}_{k}^{>}:=(\mathcal{X}_{k}\bowtie\cdots\bowtie\mathcal{X}_{d})^{\leq 1}$ for $k\in[d]$ in addition to ${\bm{\Omega}}_{\mathtt{TTS}}\mathbf{x}$ being sketched using matrices ${\bm{\Omega}}^{>}_{\mathtt{TTS},k}:=(\mathcal{G}_{k}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}\in\mathbb{F}^{PR\times n_{k}\dots n_{d}}$. We note that these matrices are themselves TTStack sketches and inherit their theoretical guarantees (Theorems 3.7 and 3.10), although they are clearly not independent of one another.
@@
Algorithm 4 Partial Contractions with TTStack
@@
2:Partial sketches $(\mathbf{W}_{1},\dots,\mathbf{W}_{d})$ with $\mathbf{W}_{k}={\bm{\Omega}}^{>}_{\mathtt{TTS},k}(\mathbf{X}^{>}_{k})^{T}\in\mathbb{F}^{Pr_{k-1}\times r_{k-1}}$.
@@ 5.2.1 Linear combinations
Explicitly assembling these block-sparse cores of size $J\chi\times n\times J\chi$ and then applying Algorithm 4 is clearly an ineffective way to compute the partial contractions. Indeed, it can be readily seen [4] that the partial sketches $\mathbf{W}_{k}={\bm{\Omega}}^{>}_{\mathtt{TTS},k}(\mathbf{S}_{k}^{>})^{T}$ for $k\geq 2$ can be formed by vertically stacking the independent partial sketches $\mathbf{W}_{k,j}={\bm{\Omega}}^{>}_{\mathtt{TTS},k}(\mathbf{S}^{>}_{k,j})^{T}$ for $j\in[J]$, obtained by applying Algorithm 4 independently to each summand TT with the sketch cores:
$\mathbf{W}_{k}=\begin{bmatrix}\alpha_{1}\mathbf{W}_{k,1}\\ \vdots\\ \alpha_{j}\mathbf{W}_{k,J}\end{bmatrix}\text{ for }k=d,\dots,2\quad\text{ and }\quad{\bm{\Omega}}_{\mathtt{TTS}}\mathbf{s}=\mathbf{w}_{1}=\sum_{j=1}^{J}\alpha_{j}\mathbf{w}_{1,j}=\sum_{j=1}^{J}\alpha_{j}{\bm{\Omega}}_{\mathtt{TTS}}\mathbf{y}_{j},$
@@ 5.2.2 Matrix-vector product
In the first step, the contraction costs $\mathcal{O}(nPRr_{H}(\chi+n))$ algebraic operations. The following one costs $\mathcal{O}(nPRr_{H}\chi(nr_{H}+R+\chi))$ algebraic operations. Proceeding recursively, the tensor contraction of ${\bm{\Omega}}_{\mathtt{TTS}}\mathbf{H}\mathbf{y}$ costs in total $\mathcal{O}(dnPRr_{H}\chi(nr_{H}+R+\chi))$ algebraic operations. This is significantly better than assembling the tensor train of ranks $R_{H}\chi$ representing $\mathbf{H}\mathbf{y}$ at a cost of $\mathcal{O}(d\chi^{2}n^{2}r_{H}^{2})$, then applying Algorithm 4 with a cost of $\mathcal{O}\bigl(dnPRr_{H}\chi(R+r_{H}\chi)\bigr)$.
(a) Tensor network representation of ${\bm{\Omega}}_{\mathtt{TTS}}\mathbf{H}\mathbf{y}$
@@
Figure 4: Graphical representation of the sketching of the matrix-vector product $\mathbf{H}\mathbf{y}$. The matrix $\mathbf{H}$ has a TT representation $(\mathcal{H}_{k})_{k\in[d]}$ of ranks $r_{H}$, $\mathbf{y}$ has a TT representation $(\mathcal{Y}_{k})_{k\in[d]}$ of ranks $\chi$, ${\bm{\Omega}}_{\mathtt{TTS}}\in\mathbb{F}^{PR\times N}$. The contractions are done from right to left, top to bottom.
@@ 5.3 Numerical Experiments
In this section, we present numerical tests of the TTStack sketch first in synthetic examples, then in the context of TT-rounding for compressing Hadamard products of tensors, and finally showcase a tensor train-adapted sketched Rayleigh-Ritz eigensolver applied to an example from quantum chemistry.
@@ 5.3.1 A Synthetic Example
We apply both Algorithms 2 and 3 with orthogonal TTStack sketches to compress a synthetic perturbed low-rank tensor $\mathcal{X}=\mathcal{X}_{s}+\varepsilon\mathcal{X}_{10}$, where $\mathcal{X}_{10}$ is a rank-10 Gaussian TT perturbation with variable noise level $\varepsilon=10^{-s}$. We consider two structures for the dominant term $\mathcal{X}_{s}$: a normalized sum of 16 rank-1 Gaussian Kronecker TT vectors, and a single rank-16 Gaussian TT vector. In both cases we fix the embedding dimension $PR=16$ and vary the block rank $R\in\{1,4,8,16\}$. By Theorem 3.10, increasing $R$ at fixed $PR$ reduces the subspace coherence constant $C_{\mathbf{Q}}(R)$, improving the injectivity parameter and, via Theorem 3.14, the rounding error bound. The two choices of $\mathcal{X}_{s}$ probe how the TT-rank structure of the dominant component interacts with this block-rank dependence.
@@ 5.3.2 Hadamard Product
Despite this representational efficiency, performing pointwise algebraic operations within the QTT format remains a significant computational bottleneck. We apply Algorithm 2 with Orthogonal TTStack sketch to compress a three-term Hadamard product, and report the accuracy and computation time as functions of the target rank.
Figure 7: Randomized rounding of Hadamard products of QTT functions. Empirical performance of Orthogonal TTStack sketches in the randomized rounding of a three-term QTT Hadamard product. The sketches are applied via Algorithm 2 to the functions defined in (11), evaluated on the discretized domain $0,1]^{3}$ over a $(2^{20})^{3}$ grid. We compare embedding dimensions $PR=r$ across varying block ranks $R\in\{1,4,8,16\}$, where markers indicate the median across 1000 independent trials, and shaded regions denote the interquartile range (25th to 75th percentiles).
@@ 5.3.3 Ground-state energy of a quantum chemistry Hamiltonian
In this experiment, we apply the Orthogonal TTStack sketch to the computation of the electronic ground-state energy of the lithium hydride (LiH) molecule in the ccpvdz basis.
The electronic Hamiltonian is written in second quantization [46], where it is naturally represented as a TTO. For the LiH molecule, the Hamiltonian corresponds to a symmetric matrix of size $2^{38}\times 2^{38}$. Starting from a Hartree-Fock initial guess $\mathcal{V}_{0}$, which has TT rank 1, we construct a Krylov basis by repeatedly applying the TTO and rounding the result via Algorithm 2 to maintain a bounded TT rank. The Orhogonal TTStack sketch plays the dual role of providing the random projections needed for randomized rounding of the new basis vector $\mathbf{b}_{j}=\mathbf{H}\mathbf{b}_{j-1}-\sum_{s<j}\alpha_{s}\mathbf{b}_{s}$ at each step, and also enabling a sketched Rayleigh-Ritz approach following [33].
@@
5:  Draw sketch ${\bm{\Omega}}_{\mathtt{TTS}}=(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}\in\ \mathbb{F}^{PR\times N}$ with parameters $(P,R)$
@@
Given the sketched basis matrix $\mathbf{C}=[{\bm{\Omega}}_{OTTS}\,\mathbf{b}_{1}\cdots{\bm{\Omega}}_{OTTS}\,\mathbf{b}_{m}]$ and the sketched action $\mathbf{D}=[{\bm{\Omega}}_{OTTS}\,\mathbf{H}\mathbf{b}_{1}\cdots{\bm{\Omega}}_{OTTS}\,\mathbf{H}\mathbf{b}_{m}]$, the Ritz matrix is obtained as the least-squares solution $\mathbf{M}=\mathbf{C}^{\dagger}\mathbf{D}$ via a QR decomposition of $\mathbf{C}$, and then diagonalized to yield the Ritz pairs and in particular an approximation $\lambda_{1}$ to the lowest eigenvalue of $\mathbf{H}$ and corresponding approximate eigenvector $\widehat{\mathbf{x}}=\sum_{j=1}^{m}y_{j}\mathbf{b}_{j}$ with $\mathbf{M}\mathbf{y}=\lambda_{1}\mathbf{y}$. In order to control the conditioning of the basis, we additionally implement a sketched Gram-Schmidt algorithm, meaning that our basis vectors are (partially) sketch-orthogonal, $({\bm{\Omega}}_{OTTS}\mathbf{b}_{i})^{*}({\bm{\Omega}}_{OTTS}\mathbf{b}_{j})\approx\delta_{ij}$, with inaccurate evaluations due to the TT-rounding steps necessary to keep ranks under control. This ensures good conditioning of the basis when ${\bm{\Omega}}_{OTTS}$ is an OSE for the (approximate) Krylov subspace. The same sketch is reused to (1) round matrix-vector products and linear combinations, (2) compute the coefficients needed to sketch-orthogonalize the Krylov basis, and (3) form the sketched problem. The resulting algorithm is summarized in Algorithm 5.
@@
Convergence results are reported on Figure 8. For each quantity of interest, namely eigenvalue error and eigenvector residual, we compare the quantities obtained from the sketched problem at each iteration, $\lambda_{1}-E_{\mathrm{ref}}$ and $\|\mathbf{D}\mathbf{y}-\lambda_{1}\mathbf{C}\mathbf{y}\|^{2}_{2}$, and the true residuals $\langle\widehat{\mathbf{x}}|\mathbf{H}|\widehat{\mathbf{x}}\rangle-E_{ref}$ and $\|\mathbf{H}\widehat{\mathbf{x}}-\lambda_{1}\widehat{\mathbf{x}}\|^{2}_{2}$. Our observation is that the approximate eigenvalue $\lambda_{1}$ is quite noisy and inaccurate, however evaluation of the true Rayleigh quotient shows a fairly steady convergence behavior, with 5 digits of accuracy obtained within 80 iterations. On the other hand, the sketched residual remains a good approximation of the true residual, as expected from the OSE / OSI guarantees of the TTStack sketch.
6 Oblivious Subspace Embedding (OSE) property for TTStack
This section begins the technical development of this paper, where we establish OSE and OSI guarantees by rigorous probabilistic techniques. To provide a pedagogical exposition, we provide full statements for the key results needed for the analysis of tensor-formatted sketch matrices such as ${\bm{\Omega}}_{\mathtt{GTT}}$ and ${\bm{\Omega}}_{\mathtt{TTS}}$.